# DITING: A Static Analyzer for Identifying Bad Partitioning Issues in TEE Applications

### v1

@@ 4. Evaluations
4.1. Environment Setup
4.1.1. Benchmark for Evaluation
Prior studies (Nong et al., 2022; Afrose et al., 2019) have created benchmarks for various static analysis tools that involve injecting vulnerabilities into code. However, none of the existing studies have created a benchmark for TEE partitioning issues. By referencing the construction methodology of the popular API-misuse dataset CryptoAPI-Bench (Afrose et al., 2019, 2023; Torres et al., 2023), we designed the benchmark PartitioningE-Bench for evaluating the abilities of DITING in bad partitioning detection. We will describe its creation steps below.
@@ 4.1.1. Benchmark for Evaluation
(a) Example code of the in-procedure cases. In this case, the critical data params may be passed to another in-TEE function instead of being used directly after entering TEE, making it more challenging to track its data-flow.
@@
Figure 7. Examples of test cases in PartitioningE-Bench.
We divide the test cases into 5 parts, including basic cases, in-procedure cases, variadic function cases, control-flow-based cases, and combined cases. These parts will cover various bad partitioning issues.
  1. (1)
    Basic cases are some simple instances described in in Section 2.2. These vulnerabilities typically occur in statements where parameters are directly used after invoking TEE.
  2. (2)
@@
As illustrated in Table 7, we totally provide 110 cases for PartitioningE-Bench, which are written in C programming language and developed for ARM TrustZone. Among these cases, 90 cases contain bad partitioning issues and 20 cases have no issues.
Table 7. The number of bad partitioning issues and no-issue cases in PartitioningE-Bench.
@@
4.1.2. Experimental Platform
DITING is implemented by Python in combination with the static analysis tool CodeQL. CodeQL is utilized to generate the data-flow for each parameter based on sources and sinks in Tables 4 to 6, and Algorithms 2 to 4 for data-flow generation. The Python code is responsible for retrieving these data-flows and identifying the locations of bad partitioning issues with the help of Algorithms 5 to 7. All experiments were conducted on a machine running Ubuntu 24.04, equipped with a 48-core 2.3GHz AMD EPYC 7643 processor, and 512GB RAM.
4.2. RQ1: Precision and Recall Evaluation
Table 8. The detection precision and recall of DITING on PartitioningE-Bench. NI is the number of partitioning issues, N is the number of detection results, and TP is the number of true positives. P, R and F1 indicates the precision, recall, and F1 score, respectively*.
@@ 4.2. RQ1: Precision and Recall Evaluation
Table 8 lists the ground truth count of bad partitioning instances in these TEE projects, alongside the number of issues reported by DITING. It can be seen that DITING achieves the highest precision of 97.06% in detecting bad partitioning of unencrypted data output. However, for the other two issues, more FPs are observed.
@@
Figure 8. Examples of FPs in DITING reports.
  1. (1)
  2. (2)
    Indexing for Input Buffer with Its Size: As shown in Fig. 8(b), although the index of array access is influenced by the input value params[0].memref.size, TEE can ensure that this input will not lead to buffer overflow. This is because that the input params[0].memref.buffer in TEE is actually a memory copy of the normal-side data, which is allocated with a size based on the above input value, and this input value is usually the size of the input data outside TEE. It means that the size of params[0].memref.buffer must not be smaller than the input value params[0].memref.size. Therefore, it is secure to directly use params[0].memref.size as an array index, even if no checks are performed on it which violates Rule 2.
  3. (3)
    Variable Assignment Using Shared Memory: Similar to the input buffer, shared memory also transmits memory address and memory size, from the normal side to TEE by params[3].memref. According to Algorithm 1, both of these parameters are treated as the type of shared memory. Meanwhile, any assignment statement containing shared memory parameters is not allowed in Rule 3. However, the variable assignment in line 1 of Fig. 8(c) can be acceptable, as it differs from the insecure shallow copy of shared memory in line 2.
According to the Recall rate in Table 8, DITING can give more comprehensive coverage of bad partitioning related to the direct use of shared memory than other issues. We also give two examples of FNs in Fig. 9 and discuss them as follow:
@@ 3 TEE_MemMove(dest, src, strlen(src));
(a) Line 15 is the FN of unencrypted data output, where params[1].memref.buffer is an output parameter.
@@
Figure 9. Examples of FNs in DITING reports.
  1. (1)
    Using Memory Operation Functions In Procedure: The destination buffer dest of the function TEE_MemMove in line 3 of Fig. 9(a) is the formal parameter of function test. Therefore, DITING recognizes the type of dest as a character array. However, as illustrated in lines 10 to 13, there are two instances of using test, one involves copying data to the in-TEE buffer str, while the other will copy data to an output. Since DITING does not identify the actual parameter type of dest as an output, it also fails to report the above second instance as bad partitioning, which violates Rule 1.
  2. (2)
    Array Access Affected by Input: Lines 3 to 5 give a for loop block, which is responsible for copying an buffer byte by byte. At this point, the input value params[2].memref.size does not directly serve as an index for array access but is instead used in the condition of for statement to constrain the value of i. Therefore, Algorithm 3 does not recognize this type of array access statement, that has broken Rule 2.
@@
4.3. RQ2: Efficiency Evaluation
Figure 10. Time cost of data-flow construction and analysis under different scales of LoC in DITING.
@@ 4.3. RQ2: Efficiency Evaluation
On our benchmark tests, we recorded the time taken by DITING to generate data-flows for three types of parameters and match these data-flows against the corresponding rules.
@@
Algorithms 2 to 4, which are responsible for constructing data-flows for parameters, involve $m$ parameters that need to be tracked and $n$ sensitive functions and data in TEE, resulting in the algorithmic complexity of them being $O(mn)$. As shown in Fig. 10, the time consumed by constructing data-flows increases from 21.53 seconds to 34.95 seconds as the lines of code (LoC) grow. Then, DITING focuses only on critical functions and data during data-flow analysis, such as TEE_MemMove and array access operations. The time spent on data-flow analysis remains relatively low and does not grow significantly with the changes of LoC (from 3.96 milliseconds to 7.13 milliseconds).
@@
Answer to RQ2: As the code scale increases from 322 LoC to 5162 LoC, the time cost of the data-flow construction for bad partitioning analysis increases from 21.53 seconds to 34.95 seconds. However, this time consumption also depends on the performance of the underlying code analysis tool. DITING still exhibits high efficiency for detecting bad partitioning issues in TEE projects (around 5.56 milliseconds).
4.4. RQ3: In-the-Wild Study
@@
We searched for TEE-related projects on GitHub using API keywords associated with TEE. After excluding TEE SDK projects, we selected the top 16 projects with the highest star counts to evaluate the performance of DITING on real-world programs, as shown in Table 9. These projects include 10 projects that have been manually analyzed in Section 2.2, along with 6 additional projects.
@@
In these projects, DITING reports 68 issues caused by bad partitioning and 55 of them are confirmed by ourselves. For example, in optee-sdp project, the log functionality implemented via snprintf copies a large amount of register data into the output parameters, which are accessible only to TEE. This compromises the confidentiality of TEE. Additionally, in projects such as Lenet5_in_OPTEE, external_rk_tee_user, there are instances where input buffers are directly used without validating their length. Attackers could exploit this bad partitioning to cause buffer overflows in TEE applications. What’s more, in secvideo project, encryption and decryption operations are directly performed on video data in shared memory, undermining the isolation provided by TEE.
@@
Table 9. The performance of DITING in detecting three kinds of bad partitioning issues on real-world TEE projects.
@@
Answer to RQ3: DITING is capable of identifying bad partitioning issues in real-world projects. Through a review of 68 issues reported by DITING in the wild, we have confirmed that 55 of them were true alarms.
5. Threats to Validity
5.1. Internal Validity
@@
The performance of DITING also heavily relies on the underlying static analysis tools. However, most tools (e.g., CodeQL and Frame-C (Correnson et al., 2024)) typically necessitate successful compilation and expert customization. This dependency limits their applicability to analyzing incomplete or uncompilable code. If these tools fail to accurately generate data-flows, the detection capabilities of DITING could be compromised. LLMDFA (Wang et al., 2024) proposed by Wang et al.makes up for this defect. Combined with the large language model, LLMDFA effectively improves the accuracy of identifying data-flow paths.
5.2. External Validity
@@
The evaluations focus on TEEs based on ARM TrustZone and Intel SGX architectures. The detection performance of DITING on other TEE architectures, such as AMD SEV (Kaplan, 2023; Mofrad et al., 2018) or RISC-V-based TEE (Nasahl et al., 2021; Suzaki et al., 2020), needs further evaluation. In addition, the evaluated projects are relatively small in scale and do not cover complex industrial-level projects. Future studies should evaluate DITING on a broader range of TEE architectures and large-scale real-world projects, thereby improving the external applicability of the experimental results.
5.3. Construct Validity
@@
Our study defines three types of rules (encrypting before output, checking value before in-TEE use, and using deep copy of shared memory) to identify bad partitioning issues. However, in practical development, there may be certain security assumptions that violate these rules without actually compromising security, potentially leading to false positives. For example, directly outputting a static string (e.g., a software version number, user id, etc.) that does not contain any sensitive data is considered secure and does not require encryption. Therefore, if these definitions are not accurate or comprehensive enough, it may lead to biased detection results. We need to update the rules based on the newly discovered bad partitioning issues in actual TEE applications to expand the detection capabilities of DITING.
6. Related Work
6.1. Vulnerabilities in TEE
@@
TEE has become integral to security-critical applications by providing isolated execution environments to protect sensitive data and computations. However, TEE applications are still affected by some vulnerabilities, as demonstrated by extensive research across various TEE technologies. Cerdeira et al. (Cerdeira et al., 2020) conducted a study on vulnerabilities in TrustZone-based systems. They have revealed critical implementation bugs, such as buffer overflows and inadequate memory protection mechanisms. Their work also underscores the role of hardware-level issues like microarchitectural side-channels in compromising TEE security. Fei et al. (Fei et al., 2021) provided a taxonomy of secure risks in Intel SGX and discussed attack vectors, such as hardware side-channel exploits and software code vulnerabilities. Similarly, Khandaker et al. (Khandaker et al., 2020) introduced the concept of COIN attacks against SGX, and they expose the insecurity of untrusted interfaces in TEE projects. Bulck et al. (Van Bulck et al., 2019) assessed the vulnerabilities at the level of the Application Binary Interface (ABI) and the Application Programming Interface (API) that can lead to memory and side-channel attacks in TEE runtime libraries.
@@
To mitigate the risk of attacks on TEE projects, some research has designed tools to detect vulnerabilities within TEE. Briongos et al. (Briongos et al., 2023) present CloneBuster for detecting cloning attacks on Intel SGX applications. CloneBuster can find whether there are same TEE application binaries running on one platform, which may roll back the TEE applications to the previous state. Ghaniyoun et al. (Ghaniyoun et al., 2023) designed TEESec, a pre-silicon framework for discovering microarchitectural vulnerabilities in TEE. TEESec can test underlying microarchitecture against data and metadata leakage across isolation boundaries.
@@
6.2. TEE Application Partitioning
Some research has attempted to mitigate vulnerabilities arising from bad partitioning by porting entire programs into TEE. For example, CryptSQLite (Wang et al., 2020) is a database system that places the entire database engine into TEE. This approach prevents attackers from stealing or tampering with critical data computed in TEE from the normal world, effectively protecting the confidentiality and integrity of data during database retrieval. RT-TEE (Wang et al., 2022) is also a cyber-physical system that puts the whole flight control code into TrustZone. This can protect the stability of equipment like drones. However, this way significantly increases Trusted Computing Base (TCB) of the applications, contradicting the principle of maintaining a small TCB for trusted code (Shinde et al., 2017; Li et al., 2019).
To address this, some other research made efforts to identify sensitive data and functions that require protection and place them inside TEE, ensuring a balance between security and efficiency. Rubinov et al. (Rubinov et al., 2016) developed an automatic approach for partitioning Android applications into critical code running in TEE and client code in the normal world. They also reduce the overhead due to transitions between the two worlds by optimizing the granularity of TEE code. Sun et al. (Sun et al., 2024) proposed dTEE that enables developers to declare tiered-sensitive variables and functions of existing IoT applications. What’s more, dTEE can automatically transform device drivers into trusted ones. Moreover, DarkneTZ (Mo et al., 2020) is a framework that limits the attack surface of the model against Deep Neural Networks with the help of Trusted Execution Environment in an edge device. They partitioned the model into sensitive layers inside TEE, and other layers executed in the normal world. This will provide reliable model privacy and defend against membership inference attacks.
The above TEE program partitioning schemes ensure software security while also maintaining the efficiency of program execution. However, as discussed in Section 2.2, improper TEE partitioning can still result in vulnerabilities. Therefore, a tool capable of analyzing and identifying bad partitioning issues in TEE implementations is essential.
7. Conclusion and Future Work

### v2

@@ IV Evaluations
IV-A Environment Setup
IV-A1 Benchmark for Evaluation
Prior studies [30, 31] have created benchmarks for various static analysis tools that involve injecting vulnerabilities into code. However, none of the existing studies have created a benchmark for TEE partitioning issues. By referencing the construction methodology of the popular API-misuse dataset CryptoAPI-Bench [31, 32, 33], we designed the benchmark PartitioningE-Bench for evaluating the abilities of DITING in bad partitioning detection. We will describe its creation steps below.
@@ IV-A1 Benchmark for Evaluation
(a) Example code of the in-procedure cases. In this case, the critical data params may be passed to another in-TEE function instead of being used directly after entering TEE, making it more challenging to track its data flow.
@@
Figure 8: Examples of test cases in PartitioningE-Bench.
We first collected code snippets with known bad partitioning issues from real-world TEE projects mentioned in Section II-C to construct the basic dataset and further expanded the dataset following the methodology in [31]. We then categorized the test cases into 5 groups, including basic cases, in-procedure cases, variadic function cases, control-flow-based cases, and combined cases. These categories cover a wide range of bad partitioning patterns and enable comprehensive evaluation of our detection approach.
  1. 1.
    Basic cases are some simple instances described in in Section II-C. These vulnerabilities typically occur in statements where parameters are directly used after invoking TEE.
  2. 2.
@@
As illustrated in Table IV, we totally provide 110 cases for PartitioningE-Bench, which are written in C programming language and developed for ARM TrustZone. Among these cases, 90 cases contain bad partitioning issues and 20 cases have no issues.
TABLE IV: The number of bad partitioning issues and no-issue cases in PartitioningE-Bench.
@@
IV-A2 Experimental Platform
DITING is implemented by Python in combination with the static analysis tool CodeQL. CodeQL is utilized to generate the data flow for each parameter based on sources and sinks in Section III-B. The Python code is responsible for retrieving these data flows and identifying the locations of bad partitioning issues with the help of Algorithms 1. All experiments were conducted on a machine running Ubuntu 24.04, equipped with a 48-core 2.3GHz AMD EPYC 7643 processor, and 512GB RAM.
IV-B RQ1: Precision and Recall Evaluation
TABLE V: The detection precision and recall of DITING on PartitioningE-Bench. NI is the number of partitioning issues, N is the number of detection results, and TP is the number of true positives. P, R and F1 indicates the precision, recall, and F1 score, respectively*.
@@ IV-B RQ1: Precision and Recall Evaluation
Table V lists the ground truth count of bad partitioning instances on PartitioningE-Bench, alongside the number of issues reported by DITING. It can be seen that DITING achieves the highest precision of 97.06% in detecting bad partitioning of unencrypted data output. However, for the other two issues, more FPs are observed.
@@
Figure 9: Examples of FPs in DITING reports.
@@ 3 TEE_MemMove(dest, src, strlen(src));
(a) Line 15 is the FN of unencrypted data output, where params[1].memref.buffer is a raw pointer passed into test.
@@
Figure 10: Examples of FNs in DITING reports.
  1. 1.
  2. 2.
    Indexing for Input Buffer with Its Size: As shown in Fig. 9(b), although the index of array access is influenced by the input value params[0].memref.size, TEE can ensure that this input will not lead to buffer overflow. This is because the input params[0].memref.buffer in TEE is actually a memory copy of the normal-side data, which is allocated with a size based on the above input value, and this input value is usually the size of the input data outside TEE. It means that the size of params[0].memref.buffer must not be smaller than the input value params[0].memref.size. Therefore, it is secure to directly use params[0].memref.size as an array index, even if no checks are performed on it which violates Rule 2.
  3. 3.
    Variable Assignment Using Shared Memory: Similar to the input buffer, shared memory also transmits memory address and memory size, from the normal side to TEE by params[3].memref, thus, both of these parameters are treated as the type of shared memory. Meanwhile, any assignment statement containing shared memory parameters is not allowed in Rule 3. However, the variable assignment in line 2 of Fig. 9(c) can be acceptable, as it differs from the insecure shallow copy of shared memory in line 4.
According to the Recall rate in Table V, DITING can give more comprehensive coverage of bad partitioning related to the direct use of shared memory than other issues. We also give two examples of FNs in Fig. 10 and discuss them as follow:
  1. 1.
    In-procedure Cases with Raw Pointers: Different from the parameter passing containing TEE_Param type in Fig. 8(a), line 15 of Fig. 10(a) passes a raw pointer (params[1].memref.buffer) to dest, a formal parameter of the function test. Due to the loss of type semantics during raw pointer passing, DITING detects the copy of sensitive data key to dest but does not flag it as a violation, as it cannot identify dest as an output parameter without its type. Additionally, in line 12, the copy of key to a local buffer str (defined within the TEE code) via the same function is considered secure. These in-procedure cases involving raw pointers are therefore difficult to detect due to the lack of type semantics.
  2. 2.
    Array Access Affected by Input: Lines 3 to 5 give a for loop block, which is responsible for copying an buffer byte by byte. At this point, the input value params[2].memref.size does not directly serve as an index for array access but is instead used in the condition of for statement to constrain the value of i. Therefore, DITING does not recognize this type of array access statement, that has broken Rule 2.
@@
IV-C RQ2: Efficiency Evaluation
@@
On our benchmark tests, we recorded the time taken by DITING to generate data flows for three types of parameters and match these data flows against the corresponding rules.
@@
As shown in Fig. 11, the time consumed by constructing data flows increases from 21.53 seconds to 34.95 seconds as the lines of code (LoC) grow. Then, DITING focuses only on critical functions and data during data flow analysis, such as TEE_MemMove and array access operations. The time spent on data flow analysis remains relatively low and does not grow significantly with the changes of LoC (from 3.96 milliseconds to 7.13 milliseconds).
@@
Answer to RQ2: As the code scale increases from 322 LoC to 5162 LoC, the time cost of the data flow construction for bad partitioning analysis increases from 21.53 seconds to 34.95 seconds. However, this time consumption also depends on the performance of the underlying code analysis tool. DITING still exhibits high efficiency for detecting bad partitioning issues in TEE projects (around 5.56 milliseconds).
Figure 11: Time cost of data flow construction and analysis under different scales of LoC in DITING.
IV-D RQ3: In-the-Wild Study
@@
Beside 10 real-world TEE projects in Section II-C, we also select another 6 projects to evaluate the performance of DITING on real-world programs, as shown in Table VI. These projects include 10 projects that have been manually analyzed in Section II-C, along with 6 additional projects.
@@
In these projects, DITING reports 154 issues caused by bad partitioning, and 130 of them are confirmed by ourselves. For example, in the optee-sdp project (an application for securely storing device data in the TEE registers), the log functionality implemented via snprintf copies a large amount of register data into the output parameters, which are accessible only to TEE. This compromises the confidentiality of TEE. Additionally, in projects such as Lenet5_in_OPTEE (a Lenet5 convolutional neural network application in ARM TrustZone) and external_rk_tee_user (a sanity testsuite for OP-TEE), there are instances where input buffers are directly used in the memory copy functions without validating their length. In particular, the PPFL project (a privacy-preserving federated learning framework with TEE) uses a large number of constants as the access indexes of the input buffer, but lacks the validation of whether the input buffer size satisfies the constant value. Therefore, PPFL has a large number of issues that violate Rule 2. Attackers could exploit the issues of input validation weaknesses to cause buffer overflows in TEE applications. Then, in the secvideo_demo project (a proof-of-concept for secure video playback on ARM TrustZone hardware), encryption and decryption operations are directly performed on video data in shared memory, undermining the isolation provided by TEE. What’s more, we have contributed five pull requests addressing those involving input validation weaknesses, which are the most feasible to fix without requiring extensive architectural changes. One of these patches [^6] https://github.com/shuaifengyun/basicAlg_use/pull/1 has already been confirmed and merged into the main branch by the maintainers.
@@
TABLE VI: The performance of DITING in detecting three kinds of bad partitioning issues on real-world TEE projects.
@@
Answer to RQ3: DITING is capable of identifying bad partitioning issues in real-world projects. Through a review of 154 issues reported by DITING in the wild, we have confirmed that 130 of them were true alarms by ourselves.
V Threats to Validity
V-A Internal Validity
@@
The performance of DITING also heavily relies on the underlying static analysis tools. However, most tools (e.g., CodeQL and Frama-C [34]) typically necessitate successful compilation and expert customization. This dependency limits their applicability to analyzing incomplete or uncompilable code. If these tools fail to accurately generate data flows, the detection capabilities of DITING could be compromised. LLMDFA [35] proposed by Wang et al.makes up for this defect. Combined with the large language model, LLMDFA effectively improves the accuracy of identifying data flow paths.
V-B External Validity
@@
The evaluations focus on TEEs based on ARM TrustZone. The detection performance of DITING on other TEE architectures, such as Intel SGX, AMD SEV [36, 37] and RISC-V-based TEE [38, 39], needs further evaluation. In addition, the evaluated projects are relatively small in scale and do not cover complex industrial-level projects. Future studies should evaluate DITING on a broader range of TEE architectures and large-scale real-world projects, thereby improving the external applicability of the experimental results.
V-C Construct Validity
@@
Our study defines three types of rules to identify bad partitioning issues. However, in practical development, there may be certain security assumptions that violate these rules without actually compromising security, potentially leading to false positives. For example, directly outputting a static string (e.g., a software version number, user id, etc.) that does not contain any sensitive data is considered secure and does not require encryption. Therefore, if these definitions are not accurate or comprehensive enough, it may lead to biased detection results. We need to update the rules based on the newly discovered bad partitioning issues in actual TEE applications to expand the detection capabilities of DITING.
VI Related Work
VI-A Vulnerabilities in TEE
@@
TEE has become integral to security-critical applications by providing isolated execution environments to protect sensitive data and computations. However, TEE applications are still affected by some vulnerabilities, as demonstrated by extensive research across various TEE technologies. Cerdeira et al. [17] conducted a study on vulnerabilities in TrustZone-based systems. They have revealed critical implementation bugs, such as buffer overflows and inadequate memory protection mechanisms. Their work also underscores the role of hardware-level issues like microarchitectural side-channels in compromising TEE security. Fei et al. [40] provided a taxonomy of secure risks in Intel SGX and discussed attack vectors, such as hardware side-channel exploits and software code vulnerabilities. Similarly, Khandaker et al. [16] introduced the concept of COIN attacks against SGX, and they expose the insecurity of untrusted interfaces in TEE projects. Bulck et al. [41] assessed the vulnerabilities at the level of the Application Binary Interface (ABI) and the Application Programming Interface (API) that can lead to memory and side-channel attacks in TEE runtime libraries.
@@
To mitigate the risk of attacks on TEE projects, some research has designed tools to detect vulnerabilities within TEE. Briongos et al. [42] present CloneBuster for detecting cloning attacks on Intel SGX applications. CloneBuster can find whether there are same TEE application binaries running on one platform, which may roll back the TEE applications to the previous state. Ghaniyoun et al. [43] designed TEESec, a pre-silicon framework for discovering microarchitectural vulnerabilities in TEE. TEESec can test an underlying microarchitecture against data and metadata leakage across isolation boundaries.
@@
VI-B TEE Application Partitioning
Some research has attempted to mitigate vulnerabilities arising from bad partitioning by porting entire programs into TEE. For example, CryptSQLite [44] is a database system that places the entire database engine into TEE. This approach prevents attackers from stealing or tampering with critical data computed in TEE from the normal world, effectively protecting the confidentiality and integrity of data during database retrieval. RT-TEE [45] is also a cyber-physical system that puts the whole flight control code into TrustZone. This can protect the stability of equipment like drones. However, this way significantly increases the Trusted Computing Base (TCB) of the applications, contradicting the principle of maintaining a small TCB for trusted code [46, 47].
To address this, some other research made efforts to identify sensitive data and functions that require protection and place them inside TEE, ensuring a balance between security and efficiency. Rubinov et al. [48] developed an automatic approach for partitioning Android applications into critical code running in TEE and client code in the normal world. They also reduce the overhead due to transitions between the two worlds by optimizing the granularity of TEE code. Sun et al. [49] proposed dTEE that enables developers to declare tiered-sensitive variables and functions of existing IoT applications. What’s more, dTEE can automatically transform device drivers into trusted ones. Moreover, DarkneTZ [50] is a framework that limits the attack surface of the model against Deep Neural Networks with the help of Trusted Execution Environment in an edge device. They partitioned the model into sensitive layers inside TEE, and other layers executed in the normal world. This will provide reliable model privacy and defend against membership inference attacks.
The above TEE program partitioning schemes ensure software security while also maintaining the efficiency of program execution. However, as discussed in Section II-C, improper TEE partitioning can still result in vulnerabilities. Therefore, a tool capable of analyzing and identifying bad partitioning issues in TEE implementations is essential.
VII Conclusion